OpenStudy (anonymous):

solve each system using substitution. 3x+2y=23 1/2x-4=y

OpenStudy (anonymous):

help please i have been stuck for hours

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

Substitute y from 2nd equation to 1st equation \[3x + 2(\frac{1}{2x}-4) = 23\] \[3x + \frac{1}{x} = 31\] \[3x^2 - 31x + 1 = 0\] Solve it then you have \[\Delta = 31^2 - 4*3 = 949\] So \[x_{1,2} = \frac{31 \pm \sqrt{949}}{6}\] You can let it be or approximate For me, I prefer to approximate results, \[x_{1} \approx 10.3 \] and \[x_{2} \approx 0.032\] Substitute then you can calculate y_1 and y_2 by yourself

OpenStudy (luigi0210):

I think you made a mistake there xD

OpenStudy (anonymous):

Now I can see you reply Luigi..

OpenStudy (anonymous):

oh wow thats a completely different way than what i was taught

OpenStudy (whpalmer4):

Yeah, it's unclear whether that is \[\frac{1}{2}x\]or\[\frac{1}{2x}\]

OpenStudy (luigi0210):

Sorry, when someone else is answering, it's polite to wait till they're done. :)

OpenStudy (anonymous):

its \[\frac{ 1 }{ 2 }x \]

OpenStudy (anonymous):

ok

OpenStudy (luigi0210):

It's good to see you again @whpalmer4 :)

OpenStudy (whpalmer4):

Well, I would do \[3x+2y=23\]\[3x+2(\frac{1}{2}x-4) = 23\]\[3x+x-8=23\]and I think you can see the rest of the way, right?

OpenStudy (whpalmer4):

@luigi0210 backatcha :-)

OpenStudy (anonymous):

i got up to there but then i got stuck

OpenStudy (luigi0210):

He's got you~

OpenStudy (whpalmer4):

Well, collect like terms...

OpenStudy (anonymous):

ok so \[4x-8=23\]

OpenStudy (whpalmer4):

Okay, add 8 to both sides

OpenStudy (anonymous):

and subtract 8 from both sides so you have 4=31 but then what it dosent work

OpenStudy (anonymous):

i mean add

OpenStudy (whpalmer4):

\[4x-8+8=23+8\]\[4x=31\]Now divide both sides by 4

OpenStudy (whpalmer4):

Yeah, you get an unattractive-looking fraction, but that's the right answer for the problem you posted...

OpenStudy (anonymous):

ok thanks

OpenStudy (whpalmer4):

then use the value you get for \(x\) to find the value of \(y\) (also una brutta figura, as it turns out)